Two weights w1 & w2 are suspended from a massless pulley . The pulley is accelerated upwards with an acceleration of g .What is the tension in the axle suspension of the pulley??

Grade:11

1 Answers

lj
6 years ago
Here is the answer to your questiong`=2gso w1`=2w1 and w2`=2w2Using T=2m1.m2.g`/(m1+m2) and substituting m1=w1/g and m2=w2/g and g` =2gWe get, tension = 4w1w2/(w1+w2)